OpenStudy (anonymous):

write an equation in point-slope for that is parallel to y=-2+3x passing through point (-2,8)

OpenStudy (anonymous):

point slope form is (y - y1) = m(x - x1) If it's supposed to be parallel to y=-2+3x, that means it has the same slope (3). So: (y - 8) = 3(x + 2)

OpenStudy (anonymous):

M=y-y1/x-x1=b/a

OpenStudy (anonymous):

the slope is (3)